so i need to do the substitution method for this equation
x-3y=8
x+y= -2
i also need to do elimination method for this equation
3x+4y= -4
5x-2y= -2

Answer by Alan3354(69443)

so i need to do the substitution method for this equation
x-3y=8
x+y= -2
---------------- Subtract
i also need to do elimination method for this equation
3x+4y= -4
5x-2y= -2
--
Multiply the 2nd eqn by 2, then add.